OpenStudy (anonymous):

Compute dz/dx for z=xln(xy)+y^3, y=cos(x^2+1)

OpenStudy (kinggeorge):

\[ -2x^2 \tan(x^2+1)-6 x \sin(x^2+1) \cos^2(x^2+1)+\log(x \cos(x^2+1))+1\]According to Wolfram Alpha